# Making Machines Sound Sarcastic: LLM-Enhanced and Retrieval-Guided Sarcastic Speech Synthesis

### v1

@@ 4.2 Sarcastic Speech Synthesis Results
Table 2 summarizes the results of different TTS systems across three perspectives: low-level acoustic metrics (MCD, pitch, energy), downstream sarcasm detection performance, and subjective naturalness ratings (Natural MOS and Sarcasm MOS). To assess downstream sarcasm detection performance, we adopt the detection architecture used in MUStARD++ with collaborative gating [11], using only the speech modality. Features extracted from synthesized audio are compared against target sarcasm labels, where higher agreement indicates stronger sarcasm expressiveness.
4.2.1 Objective Evaluation
Overall, the proposed VITS + LLaMA 3-LoRA + RAG system achieves the strongest performance across both dimensions. It obtains the lowest distortion ($9.83$ MCD) and the most stable prosodic statistics, while also yielding the highest sarcasm detection weighted F1-score (62.5%). In comparison, VITS + LLaMA 3 suffers from degraded quality (higher MCD), suggesting that raw embeddings from an untuned LLM do not align well with the synthesis task. LoRA fine-tuning mitigates this issue, and retrieval-augmented guidance further enhances both acoustic stability and sarcasm classification accuracy.
Across all systems, acoustic metrics such as MCD, pitch, and energy remain relatively stable, indicating that the improvements stem less from low-level signal fidelity and more from high-level semantic and prosodic conditioning.
4.2.2 Subjective Evaluation
In addition to objective metrics, we conducted a subjective evaluation to assess the perceptual quality and sarcastic expressiveness of the synthesized speech (Table 2). We recruited 30 listeners from diverse backgrounds who were asked to rate randomly shuffled samples from each system on two dimensions: (1) naturalness, using the standard Mean Opinion Score (MOS) on a 5-point Likert scale, and (2) sarcasm expressiveness (SMOS), reflecting how well the output conveyed sarcastic intent.
Pairwise comparison across systems reveals several important trends. First, the baseline VITS system received moderate MOS ($2.6\pm 0.1$) and SMOS ($3.2\pm 0.2$), indicating intelligible but relatively flat speech with limited sarcastic cues. When replacing the text encoder with BERT, scores remained similar ($2.5\pm 0.1$ MOS, $3.1\pm 0.2$ SMOS), suggesting that in our specific setting, BERT embeddings do not provide meaningful advantages for sarcasm conditioning.
In addition, directly integrating LLaMA 3 degraded performance ($2.0\pm 0.1$ MOS, $2.6\pm 0.2$ SMOS). Listeners frequently reported that these samples sounded less natural and had flat intonation patterns. This indicates that raw LLM embeddings are poorly aligned with intended expressive speech.
LoRA fine-tuning substantially narrowed this gap, raising naturalness to $2.6\pm 0.1$ and sarcasm expressiveness to $3.8\pm 0.2$. Compared to the baseline, listeners consistently noted clearer prosodic cues (e.g., exaggerated intonation) that aligned with sarcastic intent. This suggests that LoRA adaptation enabled LLaMA 3 embeddings to contribute pragmatically relevant sarcastic cues. Incorporating RAG provided a noticeable improvement ($2.6\pm 0.1$ MOS, $3.7\pm 0.2$ SMOS). Pairwise comparisons highlighted that RAG samples exhibited prosodic patterns more consistent with real sarcastic exemplars, making sarcastic intent more immediately recognizable.
@@ 4.2.2 Subjective Evaluation
Finally, the proposed system achieved the strongest results with $2.7\pm 0.1$ MOS and $3.8\pm 0.2$ SMOS. Taken together, the results suggest three main insights: (i) raw LLM embeddings are insufficient and may harm synthesis quality; (ii) parameter-efficient adaptation with LoRA can effectively inject pragmatic knowledge into the speech synthesis pipeline; and (iii) retrieval-augmented conditioning further refines prosodic expressiveness by grounding synthesis in real sarcastic exemplars. Together, these mechanisms enable the proposed system to generate speech that is both natural and perceptually recognizable as sarcastic.
5 Conclusion
We presented a unified framework for sarcasm-aware speech synthesis that leverages semantic conditioning from finetuned LLMs and prosodic guidance from retrieval-based exemplars. By fine-tuning LLaMA 3, our system captures semantic-level markers of sarcastic intent, while RAG provides prosodic references that enhance expressive alignment. Our results demonstrate the feasibility of modeling subtle pragmatic phenomena with TTS by combining parameter-efficient LLM adaptation and retrieval augmentation. In future work, we aim to extend this approach to other pragmatic styles such as humor, and explore cross-lingual sarcasm synthesis in diverse cultural contexts. We anticipate that this method could also serve as an effective data augmentation strategy for sarcasm detection tasks.

### v2

@@ Methods
Table 2 summarizes the evaluation of synthesized speech under different cue manipulations: semantic embeddings, prosodic exemplars, and their combination. We consider three perspectives: 1) low-level acoustic consistency, measured by mel-cepstral distortion (MCD), pitch root mean square error (RMSE), energy RMSE; 2) operationalized sarcasm expressivity, evaluated via a downstream sarcasm detection model applied to synthesized audio, and 3) subjective perception of sarcasm, assessed through two mean opinion score (MOS) tests: naturalness MOS (NMOS) and sarcasm MOS (SMOS). To assess downstream sarcasm detection performance, we adopt the detection architecture used in MUStARD++ with collaborative gating [25], using only the speech modality. Features extracted from synthesized audio are compared against target sarcasm labels, where higher agreement indicates stronger sarcasm expressiveness.
Overall, speech generated with both sarcasm-aware semantic embeddings (LoRA-fine-tuned LLaMA 3) and retrieved prosodic exemplars (RAG) exhibited the strongest performance across measures. Specifically, the combined condition achieved the lowest distortion (9.83 MCD) and the most stable prosodic statistics, while also yielding the highest downstream sarcasm detection F1-score (62.5%). Comparatively, semantic embeddings from the untuned LLaMA 3 alone resulted in higher distortion and lower detection performance, suggesting that general-purpose embeddings fail to fully capture the semantic cues critical for conveying sarcasm. Fine-tuning with LoRA improves semantic alignment, and further incorporating prosodic exemplars enhances the perceptual expressivity of the synthesized utterances. Importantly, low-level acoustic metrics (MCD, pitch, energy) remained relatively stable across conditions, indicating that improvements in sarcasm expressivity are primarily due to high-level semantic and prosodic cue manipulation, rather than changes in basic signal fidelity. Moreover, the increases observed in sarcasm detection closely mirror the subjective ratings, suggesting that the objective recognizability of sarcastic intent reliably translates to perceptual expressivity, as evaluated in the following perceptual study.
To complement the measures above, we conducted a perceptual study to evaluate how well the synthesized speech conveyed sarcasm and naturalness. We recruited 30 listeners from diverse backgrounds who were asked to rate randomly shuffled samples from each system on two dimensions: (i) perceived naturalness (NMOS) and (ii) perceived sarcasm expressivity (SMOS), both on 5-point Likert scales. Each participant listened to eight sentences generated by six systems, plus the ground truth, resulting in a total of 56 utterances. Pairwise comparison across conditions reveals several important trends. Baseline speech without semantic or prosodic guidance was generally intelligible but relatively flat, receiving moderate naturalness (2.6 ± 0.1) and sarcasm expressivity (3.2 ± 0.2). Adding general-purpose semantic embeddings (BERT) did not meaningfully improve ratings (2.5 ± 0.1 NMOS, 3.1 ± 0.2 SMOS), indicating that embeddings lacking sarcasm-specific knowledge are insufficient to enhance perceptual sarcasm cues. In addition, directly integrating LLaMA 3 degraded performance ($2.0\pm 0.1$ NMOS, $2.6\pm 0.2$ SMOS). Listeners frequently reported that these samples sounded less natural and had flat intonation patterns. This indicates that raw LLM embeddings are poorly aligned with intended expressive speech. In contrast, semantic embeddings adapted for sarcastic content significantly improved perceived sarcasm (3.8 ± 0.2 SMOS) while maintaining naturalness (2.6 ± 0.1 NMOS). Listeners consistently reported clearer prosodic patterns, such as exaggerated intonation, that aligned with the intended sarcastic meaning, suggesting that targeted semantic conditioning via LoRA adaptation can operationalize semantically and pragmatically relevant sarcastic cues in speech. Incorporating prosodic exemplars retrieved from real sarcastic utterances further enhanced listeners’ perception of sarcasm (SMOS 3.7–3.8), without reducing naturalness ($2.6\pm 0.1$ NMOS). Samples were described as having intonation and emphasis patterns more consistent with authentic sarcastic speech, making the intended sarcasm more immediately recognizable by a large portion of participants. Overall, the combined manipulation of semantic and prosodic cues produced speech that was both natural and perceptually identifiable as sarcastic ($2.7\pm 0.1$ NMOS and $3.8\pm 0.2$ SMOS). This result aligns with prior perceptual studies showing that sarcasm recognition is facilitated when semantic content and intonation are congruent [3, 32, 4]. For instance, [32] found that incongruence between context and tone led to ambiguous or delayed judgments of sarcasm. Similarly, [3, 4] demonstrated that prosodic cues systematically signal sarcastic intent in spontaneous speech, particularly when linguistic cues alone are ambiguous.
@@
Taken together, these results highlight three main key insights: (i) raw, general-purpose embeddings are insufficient and can reduce perceived expressivity; (ii) parameter-efficient adaptation with LoRA can inject semantic and pragmatic knowledge that enhances the recognition of sarcasm; and (iii) retrieval-augmented conditioning further refines prosodic expressiveness by grounding synthesis in real sarcastic exemplars. In summary, these findings suggest the complementary roles of semantic and prosodic cues in sarcasm perception.
Conclusion
This work presents a unified computational framework for modeling sarcastic speech that captures how semantic and prosodic cues interact to convey sarcastic intent. Semantic representations are derived from a LoRA-adapted LLaMA 3, encoding discourse-level cues indicative of sarcasm, while prosodic exemplars retrieved via the RAG module provide patterns of intonation and emphasis characteristic of sarcastic delivery. By integrating these components, the framework enables the study and generation of speech that reflects the subtle interplay between meaning and prosody, producing utterances that are both intelligible and perceptually recognizable as sarcastic.
Beyond improvements in synthesis quality, our findings offer insights into the cognitive mechanisms underlying sarcasm perception. Results from both objective and subjective evaluations suggest that semantic and prosodic cues make complementary contributions to the recognition of sarcasm. Crucially, neither cue alone is sufficient to reliably convey sarcastic intent, but their combination leads to the strongest perceptual effects. This supports theoretical accounts of sarcasm as a pragmatic phenomenon arising from the interaction of meaning and delivery, rather than from isolated linguistic or acoustic features.
From an applied perspective, the proposed framework demonstrates how advances in LLMs and retrieval-based conditioning can be used to model subtle pragmatic phenomena. Unlike conventional emotional TTS systems that rely on coarse affective labels, our approach offers a more flexible and cognitively grounded way of controlling expressive speech through high-level semantic cues. This has implications for the design of more socially aware conversational agents that can both recognize and produce nuanced communicative behaviors.
In future work, we plan to extend this modeling framework to other pragmatic styles such as humor, irony, and understatement, and to investigate cross-lingual sarcasm synthesis across languages and cultural contexts. In addition, we anticipate that retrieval-guided sarcastic speech synthesis could serve as an effective data augmentation strategy for downstream sarcasm detection and pragmatic inference tasks. Overall, this study highlights the value of integrating computational modeling with insights from pragmatics and speech perception to better understand and reproduce complex human communicative behaviors.